# Random Matrix Theory-guided sparse PCA for single-cell RNA-seq data

### v1

@@ Abstract
Single-cell RNA-seq provides detailed molecular snapshots of individual cells but is notoriously noisy. Variability stems from biological differences, PCR amplification bias, limited sequencing depth, and low capture efficiency, making it challenging to adapt computational pipelines to heterogeneous datasets or evolving technologies. As a result, most studies still rely on principal component analysis (PCA) for dimensionality reduction, valued for its interpretability and robustness. Here, we improve upon PCA with a Random Matrix Theory (RMT)-based approach that guides the inference of sparse principal components using existing sparse PCA algorithms. We first introduce a novel biwhitening method, inspired by the Sinkhorn–Knopp algorithm, that simultaneously stabilizes variance across genes and cells. This enables the use of an RMT-based criterion to automatically select the sparsity level, rendering sparse PCA nearly parameter-free. Our mathematically grounded approach retains the interpretability of PCA while enabling robust, hands-off inference of sparse principal components. Across seven single-cell RNA-seq technologies and four sparse PCA algorithms, we show that this method systematically improves the reconstruction of the principal subspace and consistently outperforms PCA-, autoencoder-, and diffusion-based methods in cell-type classification tasks.
I Introduction
@@
where $\bar{X}\in\mathbb{R}^{p}$ is the sample mean expression. This decomposition identifies orthogonal directions, or principal components (PCs), that capture the largest variation in the data, enabling projection of cells onto fewer axes that retain the main patterns of variability.
However, the $S$ poorly approximates the true signal contained in $\mathbb{E}[S]$ when the number of cells is comparable to the number of genes [3]. Since single-cell RNA-seq is destructive, the same cells cannot be repeatedly measured to overcome this issue, leaving only the option of sampling a very large number of cells so that $S$ accurately estimates $\mathbb{E}[S]$. In practice, this limit is rarely reached: typical experiments capture only a few thousand cells while measuring around twenty thousand genes.
For this reason, the question we address is: how can we accurately estimate $\mathbb{E}[S]$ when the number of cells $n$ and genes $p$ are large but comparable? This high-dimensional regime falls within the scope of Random Matrix Theory (RMT), and a substantial body of work leverages its results to construct estimators that minimize a chosen distance to $\mathbb{E}[S]$ [4, 5, 6, 7, 8, 9, 10]. Importantly, these estimators are rotationally invariant, meaning that the eigenvectors of $S$ remain uncorrected. In the specific context of single-cell RNA-seq, several studies have used RMT either to construct such estimators [11], or as a tool to distinguish noise from signal [12, 13, 14, 15].
In this paper, we take a different approach and focus on directly denoising the leading eigenvectors of $S$. This problem has also been studied extensively with the development of sparse PCA methods that seek sparse principal components of $S$. Although many sparse PCA algorithms have been proposed [16, 17, 18, 19, 20, 21], none have been systematically applied to realistic single-cell RNA-seq datasets. A likely reason is that sparse PCA is highly sensitive to the choice of the penalty parameter: overestimating it can introduce misleading artifacts that may easily be mistaken for biological signal. We resolve this issue by showing on a variety of single-cell RNA-seq datasets that RMT enables robust, hands-off inference of sparse principal components that better approximate the leading eigenspace of $\mathbb{E}[S]$ than standard PCs. Through systematic benchmarks on datasets with ground-truth cell type annotations, we find that sparse PCA consistently outperforms autoencoder-, diffusion-, and PCA-based methods.
Assumptions
@@
where $Y_{ij}$ are i.i.d. random variables with zero mean and unit variance, and $P=\mathbb{E}[X]$ is a low-rank matrix. The signal, also assumed low-rank, may lie in the mean $P$ (the information-plus-noise model [23]), in the covariance $B$ (the spiked separable covariance model [24, 25, 26]), or in both. We adopt standard assumptions for the separable covariance matrice model [27, 28]. Notably: the fourth moments of the entries of $Y$ are bounded, the spectral distributions of $A$ and $B$ converge to compactly supported probability distributions as $n\to\infty$ with $q=p/n$ fixed, and these limits are associated with densities $\rho_{A}$ and $\rho_{B}$, respectively. Under these assumptions, the spectral distribution of $S$ converges to a compactly supported probability distribution with density $\rho_{S}$ [27].
Methodology
@@
A central goal of RMT is to separate the contributions to the eigenspectrum of $S$ arising from the noise and from the low-rank signal. Under the assumptions above, RMT provides an analytical mapping between signal eigenvalues and the outlier eigenvalues $\lambda$ of $S$ that lie outside the support of $\rho_{S}$ [26, 28]. The eigenvectors associated with these outliers span the outlier eigenspace, which we aim to denoise with sparse PCA. Crucially, RMT not only predicts the mapping between signal and outlier eigenvalues, but also the angle between each signal eigenvector and the outlier eigenspace, and vice versa [29, 28]. This suggests that we can search for sparse PCs consistent with these angle predictions, following a maximum a posteriori inference approach [30]. In practice, however, this task is computationally demanding. We therefore simplify it by selecting the sparsity level in sparse PCA so that the inferred subspace and the outlier subspace approximately match the angle predicted by RMT.
@@
However, two obstacles arise: (i) we need estimators of $\rho_{A}$, $\rho_{B}$ to compute $\rho_{S}$ and identify the outlier eigenspace, and (ii) the RMT mapping between signal and outlier eigenspaces depends on whether the signal lies in $P$, in $B$, or in both, as detailed in App. A. To address the first obstacle, we develop an efficient algorithmic procedure called biwhitening to jointly estimate $A$ and $B$. Yet, we show that empirical estimators of $\rho_{A}$ and $\rho_{B}$ derived from these estimates fail to reconstruct the support of $\rho_{S}$, a crucial step to identify the outlier eigenspace. Lacking better estimators, we instead use our estimates of $A$ and $B$ to form the biwhitened matrix $X\leftarrow A^{-1/2}XB^{-1/2}$, for which $\rho_{S}$ is known analytically, as detailed in Section. IV. This ensures a reliable estimation of the outlier eigenspace. Moreover, this step also resolves the second obstacle: recent results show that when $A=I$, there exists a unique mapping between the signal and the outlier eigenspaces, irrespective of whether the signal lies in $P$, $B$, or both [31]. Thus, when working with biwhitened data $X\leftarrow A^{-1/2}XB^{-1/2}$, we can exploit this unique RMT mapping to guide a sparse PCA algorithm without assuming a specific model. Further discussion and illustrations of these recent findings are provided in App. A.
Contribution of this paper
For the reasons outlined above, we propose a two-step approach: (i) estimate $A$ and $B$ using a novel biwhitening algorithm, and (ii) use RMT results to guide the choice of the sparsity parameter in sparse PCA, thereby denoising the PCs of the biwhitened matrix $X\leftarrow A^{-1/2}XB^{-1/2}$. In single-cell RNA-seq analysis, PCA is most often applied to $z$-scored data to increase resolution and improve the accuracy of tasks such as cell type classification. Thus, although currently limited to biwhitened data, our approach remains particularly relevant in this setting.
@@
II Results
Estimation of $A$ and $B$
@@ II Results
Having estimated $A$ and $B$, we can construct empirical estimators for $\rho_{A}$ and $\rho_{B}$, respectively $\rho_{A}(t)\approx\sum_{i}^{n}\delta(t-c_{i}^{-2})/n$ and $\rho_{B}(t)\approx\sum_{j}^{p}\delta(t-d_{j}^{-2})/p$, and use them to solve numerically for $\rho_{S}$, as detailed in Section. IV. In Fig. 1a, we overlay this theoretical prediction with the eigenspectrum of $S$ from a single-cell RNA-seq dataset after library size normalization and $\log(x+1)$ transform. The RMT solution closely matches the data, confirming that the biwhitening procedure provides accurate estimates of $A$ and $B$. However, as shown in the inset of Fig. 1a and noted in the introduction, the RMT solution fails to correctly estimate the support of $\rho_{S}$. This failure arises because the empirical estimators of $\rho_{A}$ and $\rho_{B}$ formed from $C$ and $D$ (contain outlier eigenvalues, which are absent from the true limiting densities as $n\to\infty$ with $q=p/n$, thereby preventing the identification of the outlier eigenspace.
@@
Figure 1: Biwhitening ensures robust separation of noise and signal. Factors AA and $B$ are estimated using our Sinkhorn-Knopp Biwhitening algorithm. a. Numerical solution for the density $\rho_{S}$ from Eq. 11 and Eq. 13 using empirical estimators for $\rho_{A}$ and $\rho_{B}$. b. Eigenspectrum of the covariance matrix after biwhitening, shown alongside the Marchenko-Pastur distribution (dashed red line). The inset highlights the spectral edge, with outlier eigenvalues in pink. The close agreement with the Marchenko-Pastur prediction is supported by the small Kolmogorov-Smirnov (KS) distance and large $p$-value [35]. c. KS distance between the covariance eigenspectrum after biwhitening (red), gene-wise $z$-scoring (red), gene-wise $z$-scoring followed by cell-wise $z$-scoring (blue), and gene-wise $z$-scoring followed by cell-wise $L_{2}$ normalization (orange). All datasets were processed using $2500$ highly variable genes, following the flow charts in Fig. S2 (see Section B) For a and b, the dataset used is Luecken2021.
@@
For this reason, we work with biwhitened data, obtained using the matrices $C$ and $D$: $X\leftarrow CXD$. For such data, we expect the eigenspectrum of $S$ to closely follow the analytically known Marchenko–Pastur distribution. This is illustrated in Fig. 1b, with the inset showing a zoom on the rightmost edge of the spectrum. The Marchenko–Pastur law accurately captures the bulk support of the eigenvalues of $S$, leaving only a few eigenvalues above the spectral edge $\lambda_{+}=(1+\sqrt{q})^{2}$. Since this support is known analytically, we can directly identify the outlier eigenvalues $O$ as those exceeding $\lambda_{+}$. Their associated eigenvectors span the outlier eigenspace, equivalently represented by its orthogonal projector $W$. Finally, the quantitative agreement between the Marchenko–Pastur distribution and the observed eigenspectrum, measured by the Kolmogorov–Smirnov (KS) distance in the inset, further supports the separable covariance model as a good assumption for single-cell RNA-seq data.
@@
To illustrate the robustness of our approach, we applied it to seven datasets corresponding to seven different single-cell RNA-seq technologies [36, 37, 38, 35, 39, 40, 41]. In Fig. 1c, we compare the performance of our biwhitening algorithm against several alternative whitening strategies: (i) gene-wise $z$-scoring, (ii) gene-wise $z$-scoring followed by cell-wise $z$-scoring [12], and (iii) gene-wise $z$-scoring followed by cell-wise $L_{2}$ normalization [13]. The first method is standard processing before PCA in single-cell RNA-seq. Method (i) is the standard preprocessing step before PCA in single-cell RNA-seq. Method (ii) was proposed in [12] to better align the eigenspectrum with the Marchenko–Pastur distribution compared to (i), which we confirm in Fig. 1c. Method (iii), recently introduced as a new normalization approach [13], is in fact mathematically very similar to (ii) and produces identical performance in terms of KS distance. Across all datasets, our biwhitening approach consistently outperforms these whitening variants, with the largest gains observed when gene-wise $z$-scoring alone fails to sufficiently whiten the data, as reflected by larger KS distances.
@@
Our approach closely resembles the recently introduced BiPCA algorithm [11], which estimates biwhitening factors $C$ and $D$ by minimizing the Kolmogorov–Smirnov distance between the eigenspectrum and the Marchenko–Pastur distribution. However, BiPCA is restricted to operating at the level of raw counts, whereas our method can be applied at multiple stages of preprocessing: directly on counts, after library-size correction, or even after log-normalization as shown in Fig. 1. In Fig. S6a, we show that our biwhitening algorithm performs almost identically to BiPCA when applied to count data. In particular, Fig. S6a,b demonstrate that it recovers biwhitening factors $C$ and $D$ nearly identical to those of BiPCA, and this irrespective of the number of highly variable genes selected. This highlights that, at least on counts, the factors obtained with Alg. 1 are near-minimizers of the Kolmogorov–Smirnov distance between the eigenspectrum of $S$ and the Marchenko–Pastur distribution.
RMT-guided sparse PCA

### v2

@@ Abstract
Single-cell RNA-seq provides detailed molecular snapshots of individual cells but is notoriously noisy. Variability stems from biological differences and technical factors, such as amplification bias and limited RNA capture efficiency, making it challenging to adapt computational pipelines to heterogeneous datasets or evolving technologies. As a result, most studies still rely on principal component analysis (PCA) for dimensionality reduction, valued for its interpretability and robustness, in spite of its known bias in high dimensions. Here, we improve upon PCA with a Random Matrix Theory (RMT)-based approach that guides the inference of sparse principal components using existing sparse PCA algorithms. We first introduce a novel biwhitening algorithm which self-consistently estimates the magnitude of transcriptomic noise affecting each gene in individual cells, without assuming a specific noise distribution. This enables the use of an RMT-based criterion to automatically select the sparsity level, rendering sparse PCA nearly parameter-free. Our mathematically grounded approach retains the interpretability of PCA while enabling robust, hands-off inference of sparse principal components. Across seven single-cell RNA-seq technologies and four sparse PCA algorithms, we show that this method systematically improves the reconstruction of the principal subspace and consistently outperforms PCA-, autoencoder-, and diffusion-based methods in cell-type classification tasks.
1 Introduction
@@
where $\bar{X}\in\mathbb{R}^{p}$ is the sample mean expression. This decomposition identifies orthogonal directions, or principal components (PCs), that capture the largest variation in the data, enabling projection of cells onto fewer axes that retain the main patterns of variability. This approach is the primary dimensionality reduction method used in popular single-cell RNA-seq analysis packages such as Scanpy [50] and Seurat [46].
When many more cells are sampled than genes are measured, $S$ converges to $\mathbb{E}[S]$, and its eigenspectrum and eigenvectors reliably estimate those of $\mathbb{E}[S]$. In typical single-cell RNA-seq experiments, however, the number of cells is comparable to the number of genes. In this high-dimensional regime, the leading principal components of $S$ are poor estimators of those of $\mathbb{E}[S]$. In practice, the extent of this error is quantified by computing the overlap between the two principal subspaces. This overlap is maximal when $p/n\rightarrow 0$ and decreases toward zero as $p/n$ increases. Because single-cell RNA-seq is destructive, the same cells cannot be repeatedly measured to average out this variability, leaving accurate estimation of the principal components of $\mathbb{E}[S]$ a central challenge.
For this reason, the question we address is: how can we accurately estimate the PCs of $\mathbb{E}[S]$ when the number of cells $n$ and genes $p$ are large but comparable? This high-dimensional regime falls within the scope of Random Matrix Theory (RMT), which provides tools to predict the eigenspectrum of $S$, as well as the overlap between the leading eigenspaces of $S$ and $\mathbb{E}[S]$. A substantial body of work leverages these results to construct better estimators of $\mathbb{E}[S]$, with most approaches correcting leading eigenvalues of $S$ so as to minimize a chosen distance to $\mathbb{E}[S]$ [12, 29, 41, 18, 19, 30, 11]. Importantly, these improved estimators are rotationally invariant, meaning that the eigenvectors of $S$ remain uncorrected. In the specific context of single-cell RNA-seq, several studies have used RMT either to construct such estimators [45] or as a tool to distinguish noise from signal [39, 27, 3, 31].
In this paper, we take a different approach and focus on directly denoising the principal components of $S$. This problem has also been studied extensively through sparse PCA methods, which seek sparse principal components of $S$. To do so, these approaches augment PCA with sparsity constraints imposed on the principal components. Many sparse PCA algorithms have been proposed [26, 13, 25, 54, 14, 22], each implementing a variation of the same idea. However, none has been systematically applied to realistic single-cell RNA-seq datasets. A likely reason is that sparse PCA is highly sensitive to the choice of the penalty parameter: overestimating it can introduce misleading artifacts that may be mistaken for biological signal. We resolve this issue by showing, on a variety of single-cell RNA-seq datasets, that RMT enables robust, hands-off inference of sparse principal components that better approximate the leading eigenspace of $\mathbb{E}[S]$ than standard PCs. Through systematic benchmarks on datasets with ground-truth cell type annotations, we find that sparse PCA consistently outperforms autoencoder-, diffusion-, and PCA-based methods on the task of cell type annotation.
2 Results
2.1 Assumptions and methodology
@@
where $Y_{ij}$ are i.i.d. random variables with zero mean and unit variance, and $P=\mathbb{E}[X]$ is a low-rank matrix. The signal, also assumed low-rank, may lie in the mean $P$ (the information-plus-noise model [8]), in the covariance $B$ (the spiked separable covariance model [24, 5, 51]), or in both. While this central assumption is largely supported by recent findings on various single-cell RNA-seq datasets [28, 45], we will also confirm its validity on the data used in this paper. We adopt standard assumptions for the separable covariance matrice model [15, 16]. Notably: the fourth moments of the entries of $Y$ are bounded, the spectral distributions of $A$ and $B$ converge to compactly supported probability distributions as $n\to\infty$ with $q=p/n$ fixed, and these limits are associated with densities $\rho_{A}$ and $\rho_{B}$, respectively. Under these assumptions, the spectral distribution of $S$ converges to a compactly supported probability distribution with density $\rho_{S}$ [15].
@@
A central goal of RMT is to separate the contributions to the eigenspectrum of $S$ arising from the noise and from the low-rank signal. Under the assumptions above, RMT provides an analytical mapping between signal eigenvalues and the eigenvalues $\lambda$ of $S$ that lie outside the support of $\rho_{S}$ [51, 16]. The eigenvectors associated with these outlier eigenvalues span the outlier eigenspace, which we aim to denoise with sparse PCA. Crucially, RMT not only predicts the mapping between signal and outlier eigenvalues, but also the angle between each signal eigenvector and the outlier eigenspace, and vice versa [9, 16]. This suggests that we can search for sparse PCs consistent with these angle predictions, following a maximum a posteriori inference approach [40]. In practice, however, this task is computationally demanding. We simplify it by selecting the sparsity level in sparse PCA so that the inferred subspace and the outlier subspace approximately match the angle predicted by RMT.
@@
However, two obstacles arise: (i) we need estimators of $\rho_{A}$ and $\rho_{B}$ to compute $\rho_{S}$ and identify the outlier eigenspace, and (ii) the RMT mapping between signal and outlier eigenspaces depends on whether the signal lies in $P$, in $B$, or in both, as detailed in SI Appendix A. To address the first obstacle, building on recent advances in matrix biwhitening [45], we develop a novel algorithm to jointly estimate $A$ and $B$ without requiring assumptions on the noise distribution. However, we show that empirical estimators of $\rho_{A}$ and $\rho_{B}$ derived from these estimates fail to reconstruct the support of $\rho_{S}$, a crucial step to identify the outlier eigenspace. Lacking better estimators, we instead use our estimates of $A$ and $B$ to form the biwhitened matrix $X_{\mathrm{bw}}=A^{-1/2}XB^{-1/2}$, for which $\rho_{S}$ is known analytically, as detailed in Section 4. This ensures reliable estimation of the outlier eigenspace for the biwhitened matrix.
Moreover, this biwhitening step also resolves the second obstacle: recent results show that for left-whitened data $X_{\mathrm{lw}}=A^{-1/2}X$, there exists a unique mapping between the signal and outlier eigenspaces, irrespective of whether the signal lies in $P$, $B$, or both [33]. This result carries over to biwhitened data $X_{\mathrm{bw}}$, meaning that by working with $X_{\mathrm{bw}}$, we can guide a sparse PCA algorithm without assuming a specific model for where the signal lies. Further discussion and illustrations of these findings are provided in SI Appendix A. For these reasons, we propose a two-step approach: (i) estimate $A$ and $B$ using our novel biwhitening algorithm, and (ii) use RMT results to guide the choice of the sparsity parameter in sparse PCA and denoise the PCs of the biwhitened matrix $X_{\mathrm{bw}}$.
@@
2.2 A novel biwhitening algorithm ensures a robust separation of signal and noise
@@
Having estimated $A$ and $B$, we can construct empirical estimators for $\rho_{A}$ and $\rho_{B}$, respectively $\rho_{A}(t)\approx\sum_{i}^{n}\delta(t-c_{i}^{-2})/n$ and $\rho_{B}(t)\approx\sum_{j}^{p}\delta(t-d_{j}^{-2})/p$, and use them to solve numerically for $\rho_{S}$, as detailed in Section. 4. In Fig. 1a, we overlay this theoretical prediction with the eigenspectrum of $S$ from a single-cell RNA-seq dataset after library size normalization and $\log(x+1)$ transform. The RMT solution closely matches the data, confirming that the biwhitening procedure provides satisfying estimates of $A$ and $B$. However, as shown in the inset of Fig. 1a and noted in the introduction, the RMT solution fails to correctly estimate the support of $\rho_{S}$. This failure arises because the empirical estimators of $\rho_{A}$ and $\rho_{B}$ formed from $C$ and $D$ contain outlier eigenvalues, which are absent from the true limiting densities as $n\to\infty$ with $q=p/n$, thereby preventing the identification of the outlier eigenspace.
@@
Figure 1: Validity of the separable covariance model. The factors $A$ and $B$ are estimated using our Sinkhorn-Knopp Biwhitening algorithm. a. Numerical solution for the density $\rho_{S}$ from Eq. 11 and Eq. 13 using empirical estimators for $\rho_{A}$ and $\rho_{B}$. b. Eigenspectrum of the covariance matrix after biwhitening, shown alongside the Marchenko-Pastur distribution (dashed red line). The inset highlights the spectral edge, with outlier eigenvalues in pink. The close agreement with the Marchenko-Pastur prediction is supported by the small Kolmogorov-Smirnov (KS) distance and large $p$-value [36]. c. KS distance between the covariance eigenspectrum after biwhitening (red), gene-wise $z$-scoring (red), gene-wise $z$-scoring followed by cell-wise $z$-scoring (blue), and gene-wise $z$-scoring followed by cell-wise $L_{2}$ normalization (orange). All datasets were processed using $2500$ highly variable genes, following the flow charts in Fig. S2 (see SI Appendix B) For a and b, the dataset used is Luecken2021.
@@
For this reason, we work with biwhitened data, obtained using the matrices $C$ and $D$: $X_{\mathrm{bw}}=CXD$. For such data, we expect the eigenspectrum of $S$ to closely follow the analytically known Marchenko–Pastur distribution. This is illustrated in Fig. 1b, with the inset showing a zoom on the rightmost edge of the spectrum. The Marchenko–Pastur law accurately captures the bulk support of the eigenvalues of $S$, leaving only a few eigenvalues above the spectral edge $\lambda_{+}=(1+\sqrt{q})^{2}$. Since this support is known analytically, we can directly identify the outlier eigenvalues $O$ as those exceeding $\lambda_{+}$. Their associated eigenvectors span the outlier eigenspace, equivalently represented by its orthogonal projector $W$. Finally, the agreement between the Marchenko–Pastur distribution and the observed eigenspectrum, quantified by the Kolmogorov–Smirnov (KS) distance and its associated p-value in the inset, confirms that the separable covariance model is a statistically valid assumption for single-cell RNA-seq data. From a modeling perspective, this means that, with the number of genes and cells used in this example, correlation structure beyond a diagonal and separable covariance model is not statistically significant. Importantly, this does not rule out the existence of such correlations, which could be revealed by analyzing more cells than in Fig. 1.
@@
To illustrate the robustness of our approach, we applied it to seven datasets corresponding to seven different single-cell RNA-seq technologies [20, 53, 21, 36, 38, 46, 49] (see Section 4). In Fig. 1c, we compare the performance of our biwhitening algorithm against several alternative whitening strategies: (i) gene-wise $z$-scoring, (ii) gene-wise $z$-scoring followed by cell-wise $z$-scoring [39], and (iii) gene-wise $z$-scoring followed by cell-wise $L_{2}$ normalization [27]. Method (i) is the standard preprocessing step before PCA with single-cell RNA-seq data. Method (ii) was proposed in [39] to better align the eigenspectrum with the Marchenko–Pastur distribution compared to (i), which we confirm in Fig. 1c. Method (iii), recently introduced as a new normalization approach [27], is in fact mathematically very similar to (ii) and produces identical performance in terms of KS distance. Across all datasets, our biwhitening approach consistently outperforms these whitening variants, with the largest gains observed when gene-wise $z$-scoring alone fails to sufficiently whiten the data, as reflected by larger KS distances.
@@
Our approach closely resembles the recently introduced BiPCA algorithm [45], which also uses the Sinkhorn-Knopp algorithm to estimate biwhitening factors $C$ and $D$. However, BiPCA can only operate on data for which the variance of gene expression is quadratically related to its mean. This assumption holds at the level of counts, but breaks down at later stages of data processing. Instead, our method self-consistently estimates the variance of gene expression without assuming any specific relationship with the mean expression. This modification is the main novelty of our approach, as it allows us to biwhiten the data at any stage of preprocessing: directly on counts, after library-size correction, or even after log-normalization, as shown in Fig. 1. In Fig. S6a, we show that our biwhitening algorithm performs almost identically to BiPCA when applied to count data. In particular, Fig. S6a,b demonstrate that it recovers biwhitening factors $C$ and $D$ that are nearly identical to those of BiPCA, irrespective of the number of highly variable genes selected.
2.3 RMT-guided sparse PCA estimates the signal eigenspace better than standard PCA